# Brownian motion of a single microparticle partially engulfed by a giant lipid vesicle

### v1

@@ II.2.1 Experiments
In this section, we aim to describe the equilibrium dynamics of particle-vesicle systems. A distribution of tension and adhesion energy allowed us to observe various penetration depths, described by a dimensionless parameter $d/a$, where $a$ corresponds to the particle radius and $d$ the distance between the immersed pole of the particle and the membrane surface, as shown in Fig. 2A. Particles exhibit a pinning of the contact line and show anisotropic motion constrained by the curvature of the GUV. Therefore, it is necessary to project the experimentally observed 2D path into a 3D trajectory to accurately describe the diffusion of the particle on the vesicle surface. This trajectory reconstruction assumes that the vesicle has a spherical shape, which is an approximation given that gravitational effects and adhesion to the substrate may cause a flattening of the vesicle. Nevertheless, this approximation is quite suitable as long as the particle does not stray too far from the focal plane of observation where the radius of curvature is measured, which is the case for our system. Fig. 2B represents the 3D reconstruction of the diffusion of a partially engulfed particle, based on its apparent 2D trajectory. The mean squared displacement was calculated using the method of Shigyou et al. (see section V.8) and examples of MSD corresponding to free, fully and half engulfed particles are plotted in Fig. 2C. In addition to the various degrees of engulfment observed experimentally, some vesicles may also contain substructures such as lipid aggregates, filaments, smaller vesicles, or may show some multilamellar regions, which could increase the friction experienced by the particle. Fig. 3 summarizes these effects and depicts the evolution of translational friction between a silica particle and a GUV containing 5 molDOTAP. Yellow circles correspond to particles engulfed by vesicles without substructures observable by fluorescence microscopy, while pink circles correspond to cases where substructures are present (see Fig. 2A for $d/a=-1$, the vesicle exhibits substructures near the particle). The dashed and dotted blue curves are associated to numerical calculations by Fischer et al. describing the friction of a sphere passing through an incompressible two-dimensional liquid of viscosity $\eta_{m}$ (membrane), separating two liquids of infinite depth and viscosity $\eta_{w}$ (water) [13]. For $-2<d/a<0$ the friction force associated with the translational motion of the sphere at velocity $U$ is defined by:
@@
This model assumes no rotation for a particle in contact with the membrane and no deformation of the interface, remaining flat for every degree of the particle penetration (see inset Fig. 3). Dissipation results from the bulk viscosity seen by the spherical particle showing a size significantly larger than the membrane thickness and by the viscosity of the flat membrane, which forms a hole to accommodate the particle inclusion. Therefore, the cases $d/a=-2$ and $d/a=0$ are identical and the friction follows a parabolic form. However, in an endocytosis-like process, a particle does not perforate the bilayer but induces deformation, forcing the membrane to locally match the particle shape. This symmetry breaking is not considered in current models and may explain the deviations observed between our experimental data and the numerical calculations. Note that although a discrepancy exists, experimental results can be reasonably well fitted by the hydrodynamic model [13] with a viscosity $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$, both showing a maximum friction when the spherical particle is symmetrically immersed, with its equator coinciding with the membrane plane. However, some data cannot be fitted with this value of $\eta_{m}$ and correspond to particles interacting with vesicles containing substructures, which may increase the effective two dimensional membrane viscosity. For fluid bilayers, the molecular rearrangement of lipids prevents the generation of shear stress across the membrane, thus the latter can be described as a thin region of liquid with viscosity $\eta$ and thickness $\delta$, without generation of velocity gradient, or equivalently, can be treated as an anisotropic material with a constant surface shear viscosity $\eta_{m}=\delta\eta$. Therefore, the increase in friction between particles and GUVs containing substructures in their vicinity can be attributed to the linear dependence of surface shear viscosity on membrane thickness, and hence on the number of lipid monolayers. Consequently, the two highest values of the friction (pink circles), deviating from the trend ($\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$), can be described by a (2.5 times) higher effective viscosity, possibly indicating the presence of a second bilayer.
@@ II.2.2 Simulations
Several studies have demonstrated the existence of a variable water gap thickness in solid surface-membrane systems. For instance, Rädler et al. [21] investigated the weak interaction of giant vesicles with flat rigid substrates and found a mean separation distance of about 40 nm. Other studies have reported the existence of a 1-3 nm water gap in supported bilayer systems [22] [23] (given the strong adhesion in our system, the particle-membrane equilibrium distance is likely near the lower bound of this range). Concerning viscosity, several authors have measured values for DOPC bilayers ranging from 2 to 15 $\cdot 10^{-9}\ \mathrm{Pa.s.m}$ [24] [25] [26]. Fig. 5 shows the calculated values of the translational friction coefficient as a function of penetration depth for different water gap thicknesses and membrane viscosities, reflecting the range of data obtained in these different studies, with a water gap of 5 – 10 nm and viscosity ranging from 2 to 10 $\cdot 10^{-9}\ \mathrm{Pa.s.m}$. Our simulations exhibit a trend similar to Fischer et al. calculations, with maximum friction at half-engulfment. However here, the deformed shape of the membrane induced by particle wrapping is taken into account and the evolution of friction is no longer symmetrical with respect to the position $d/a=-1$. These simulations provide a correction to the hydrodynamic model [13], allowing a more accurate description of our experimental results, also displaying asymmetry. In our simulations, the presence of a water gap separating the particle and the membrane was also taken into account. A reduction in the thickness of this water layer results in increased friction, more pronounced when the membrane is highly viscous. This variation is minor for a viscosity $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$ but becomes significant for a viscosity $\eta_{m}=10^{-8}\ \mathrm{Pa.s.m}$. In our experiments, with a lipid bilayer viscosity estimated at $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$, the thickness of the water gap has thus a small effect on friction. Hence, dissipation primarily occurs in the regions close to the junction between the unbound membrane plane and the cap (formed by the particle portion engulfed and covered by the membrane, see Fig. 4A).
Figure 5: Translational friction of a spherical particle (radius $a=0.6\,\mathrm{\mu m}$) as a function of penetration depth through a membrane, modeled as a layer of incompressible fluid (width $\delta=5\,\mathrm{nm}$) and constant surface shear viscosity $\eta_{m}$, separated by two identical regions of incompressible liquid (water). A water gap (thickness $l$) between the particle and the membrane is considered in the simulations. Our numerical results are compared to Fischer et al. calculations for 2D viscosites $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$ and $\eta_{m}=10^{-8}\ \mathrm{Pa.s.m}$ (blue dotted and dashed curves). The points from the simulation, represented as circles and diamonds, are connected by a solid curve serving as a visual guide and does not represent an interpolation.
@@
Figure 6: (A) Schematic representations and bright-field images of a particle wrapped by a vesicle bilayer and a particle at the interface between the internal aqueous phase and the external oil phase of a droplet, stabilized by a lipid monolayer. The lipid composition used for the fabrication of the vesicles and droplets is identical. (B) Typical two-dimensional trajectories associated with the two configurations, for an acquisition time of 10 s and vesicle/droplet radii ranging from 3.5 to 7 µm. (C) Mean squared displacement as a function of lag time for particles wrapped by a lipid bilayer, with different penetration depths and (D) particles wetting a lipid monolayer. In the case of particles encapsulated in droplets, the MSD scales linearly with time, even for longer timescales ($\Delta t\gg 0.03$ s), characteristic of purely Brownian motion, but exhibits a lower diffusion coefficient compared to particles wrapped by a bilayer (see insets). In this latter case, the transport is hindered and manifests as a sublinear increase of the MSD at longer timescales ($\mathrm{MSD}\propto\Delta t^{\alpha}$, with $\alpha=0.78\pm 0.14$ based on all the data analyzed). The diffusion is systematically lower in both cases compared to the Stokes-Einstein predictions for a sphere in the bulk (red curve).
@@ III Particle diffusion at the water/lipid monolayer/oil interface
The analysis of the diffusion of particles wrapped by GUVs, potentially containing substructures near their contact zone, indicates an influence of the effective membrane thickness on translational friction. To investigate further the effect of the layer thickness, we performed experiments using lipid monolayers showing a thickness that is half that of lipid bilayers. We encapsulated particles in water droplets stabilized by a lipid monolayer (same composition as the vesicles) and dispersed in oil (dodecane, $\eta_{o}=1.324\cdot 10^{-3}\ \mathrm{Pa.s}$ [27]). Unlike GUVs, where particles spontaneously adhere to the outer leaflet of the lipid bilayer during the observation phase, here, the particles are directly encapsulated during the fabrication process. Fig. 6A illustrates the two systems: at the bottom, a particle partially engulfed by a vesicle bilayer and at the top, a particle encapsulated in a droplet, in close proximity with the monolayer. In pure water, the particle does not adhere to the interface and diffuses freely in three dimensions within the aqueous droplet. The addition of NaCl in the aqueous phase (10 mM) results in a reduction of the Debye length and causes the particle to approach the interface. In this configuration, the particle does not deform the lipid monolayer but is in a state of partial wetting, moving along the curvature of the droplet and exploring a larger region than particles wrapped by a GUV bilayer. Fig. 6B depicts the trajectories of particles encapsulated in droplets and wetting the interface (green) and of particles partially engulfed by GUVs (purple), over the same acquisition time (10 s), for vesicle/droplet radii ranging from 3.5 to 7 µm, clearly showing this phenomenon. Although the trajectories of particles encapsulated in droplets are more extensive, the mean squared displacement evaluated at short times (MSD linear with the lag time for $\Delta t<0.03$ s) reveals slower diffusion compared to particles wrapped by lipid bilayers. MSD analysis for particle-bilayer systems at longer times reveals a sublinear behaviour characteristic of confined diffusion dynamics (see Fig. 6C and D). The crossing from the diffusive and subdiffusive regimes occurs around $\Delta t\simeq 0.05$ s, and the subdiffusion is characterized with an exponent $\alpha=0.78\pm 0.14$ (no correlation was observed between $d/a$ and $\alpha$). This phenomenon is ubiquitous in cellular biology and has also been studied using model systems. Several works have shown that the presence of substructures (buds) [24] and membrane crowding induced by proteins [28] or lipid domains [29] lead to anomalous transport. In our system, the vesicles are made of only two types of lipids in the liquid phase (and 1% of a fluorescently labeled lipid), and thus diffusion cannot be perturbed by complex systems as proteins. The presence of substructures may explain this type of long time diffusion, but in our system, it appears even in GUVs without optically visible substructures. The finite-size effects induced by the spherical shape of the vesicle are not the cause of this confinement either, as it does not occur for a particle encapsulated in a droplet of the same size. Another hypothesis explaining particle subdiffusion could be the formation of an electrostatically induced DOTAP-rich zone in its vincinity. Among the various factors possibly contributing to particle subdiffusion, the role of membrane elasticity has also been investigated. Daddi-Moussa-Ider et al. analytically demonstrated that the motion of a particle near a membrane causes its deformation, followed by relaxation, generating a flow acting back on the particle with a certain delay. This memory effect, induced by the elastic nature of the membranes, is responsible for a transient subdiffusive phenomenon [30]. Although their work describes the diffusion of a particle near a membrane rather than being wrapped, this interpretation remains compelling in explaining the experimentally observed anomalous diffusion.
@@
Going back to the short lag time diffusive regime, we measured a particle diffusion at the lipid monolayer interface significantly lower than for lipid bilayers even if the thickness of the monolayer is half that of the bilayer. Using the hydrodynamic model [13] to fit our experimental results for monolayers yields a 2D viscosity of approximately $1.5\cdot 10^{-8}\ \mathrm{Pa.s.m}$, which is substantially higher than the estimated viscosity of lipid bilayers, even those containing substructures (in the range $\eta_{m}=2\cdot 10^{-9}-5\cdot 10^{-9}\ \mathrm{Pa.s.m}$, see Fig. 3). In this model and our simulations, the interface is described as flat (no roughness), since interfacial tension stresses dominate over viscous stresses, with a capillary number $Ca=\eta U/\sigma_{wo}\ll 1$ where $U$ is the particle velocity and $\sigma_{wo}$ is the water-oil interfacial tension. The triple-phase contact line between the particle and the two immiscible liquids appears stationary at the microscopic scale, but at the molecular level, it undergoes significant thermal motion, leading to additional energy dissipation. To more precisely capture this behavior, we used the approach of Boniello et al. [31] where a spherical microparticle diffuses at the water/air interface with a contact angle $\theta$ (related to the penetration depth by $1+\cos\theta=-d/a$), satisfying the local equilibrium of the triple contact line. Thermally activated displacements of the interface at the contact line induce additional random forces acting on the particle, which are associated with an additional viscous friction through the fluctuation-dissipation theorem. The triple-phase contact line fluctuates around its average position with a characteristic deformation length scale denoted as $\lambda$, inducing a line friction acting on the microparticle surface $\gamma_{tr,L}$ defined by:
@@
with $\chi=\arctan({4\sin\theta/\pi})$ being the angle between the horizontal plane and the tangents of the interface on the particle surface and $\tau=\kappa_{0}^{-1}$ corresponding to the correlation time of the interfacial fluctuations. This correlation time differs from that applied in the article by Boniello et al. describing a water/air interface. Here, both media are liquids with viscosities $\eta_{w}$ (water) and $\eta_{o}$ (oil) and the equilibrium frequency of random molecular displacements, $\kappa^{0}$, is defined by the following relation [32]:
@@
where $v_{w}=0.03\ \mathrm{nm}^{3}$ and $v_{o}=0.38\ \mathrm{nm}^{3}$ denote the molecular volume of water and oil, $h$ the Planck constant and $\sigma_{wo}(1+\cos\theta)$ corresponds to the reversible work of adhesion between the solid and the liquid, while considering the presence of the other liquid phase. Fig. 7 schematically represents the particle diffusing at the water/lipid monolayer/oil interface and the various parameters involved. The total friction experienced by the particle diffusing at the lipid monolayer interface is equal to the sum of two uncorrelated terms $\gamma_{tr}=\gamma_{tr,H}+\gamma_{tr,L}$. For the calculation of the hydrodynamic friction $\gamma_{tr,H}$, we approximate the system as a particle dragged along a liquid-liquid interface, with both liquids having identical viscosities, using the result of the hydrodynamic model in [13], the ratio $\eta_{o}/\eta_{w}=1.324$ being close to 1. We set the surface viscosity of the monolayer to $\eta_{m}/2$, where $\eta_{m}$ is the viscosity of a membrane bilayer with the same lipid composition ($\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$, see Fig. 3). The interfacial tension for similar systems was measured using the pendant drop technique and found to vary between $2\text{-}5\,\mathrm{mN.m^{-1}}$ [33] [34]. Considering a small contact angle $5^{\circ}<\theta<10^{\circ}$, the characteristic deformation length $\lambda$ ranges between 2.7 and 4.8 nm which can be related to thermally excited capillary waves for low tension interfaces [35] and is comparable to the length of a lipid molecule. This analysis reveals a fundamental difference between the diffusion mechanisms of a particle wrapped by a lipid bilayer, well described by a hydrodynamic model, and a particle diffusing near the water/lipid monolayer/oil interface of a droplet. For the latter system, fluctuations of the three-phase contact line must be taken into account, as they introduce an additional contribution to the translational drag experienced by the particle diffusing at the interface.

### v2

@@ II.2.1 Experiments
In this section, we aim to describe the equilibrium dynamics of particle-vesicle systems. A distribution of tension and adhesion energy allowed us to observe various penetration depths, described by a dimensionless parameter $d/a$, where $a$ corresponds to the particle radius and $d$ the distance between the immersed pole of the particle and the membrane surface, as shown in Fig. 3A.
Particles exhibit a pinning of the contact line and show anisotropic motion constrained by the curvature of the GUV. Therefore, it is necessary to project the experimentally observed 2D path into a 3D trajectory to accurately describe the diffusion of the particle on the vesicle surface. This trajectory reconstruction assumes that the vesicle has a spherical shape, which is an approximation given that gravitational effects and adhesion to the substrate may cause a flattening of the vesicle. Nevertheless, this approximation is quite suitable as long as the particle does not stray too far from the focal plane of observation where the radius of curvature is measured, which is the case for our system. Fig. 3B represents the 3D reconstruction of the diffusion of a partially engulfed particle, based on its apparent 2D trajectory. The mean squared displacement was calculated using the method of Shigyou et al. (see section V.8) and examples of MSD corresponding to free, fully and half engulfed particles are plotted in Fig. 3C.
In addition to the various degrees of engulfment observed experimentally, some vesicles may also contain substructures such as lipid aggregates, filaments, smaller vesicles, or may show some multilamellar regions, which could increase the friction experienced by the particle. Fig. 4 summarizes these effects and depicts the evolution of translational friction between a silica particle and a GUV containing 5 molDOTAP. Yellow circles correspond to particles engulfed by vesicles without substructures observable by fluorescence microscopy, while pink circles correspond to cases where substructures are present (see Fig. 3A for $d/a=-1$, the vesicle exhibits substructures near the particle). The dashed and dotted blue curves are associated to numerical calculations by Fischer et al. describing the friction of a sphere passing through an incompressible two-dimensional liquid of viscosity $\eta_{m}$ (membrane), separating two liquids of infinite depth and viscosity $\eta_{w}$ (water) [20]. For $-2<d/a<0$ the friction force associated with the translational motion of the sphere at velocity $U$ is defined by:
@@
This model assumes no rotation for a particle in contact with the membrane and no deformation of the interface, remaining flat for every degree of the particle penetration (see inset Fig. 4). Dissipation results from the bulk viscosity seen by the spherical particle showing a size significantly larger than the membrane thickness and by the viscosity of the flat membrane, which forms a hole to accommodate the particle inclusion. Therefore, the cases $d/a=-2$ and $d/a=0$ are identical and the friction follows a parabolic form. However, in an endocytosis-like process, a particle does not perforate the bilayer but induces deformation, forcing the membrane to locally match the particle shape. This symmetry breaking is not considered in current models and may explain the deviations observed between our experimental data and the numerical calculations. Note that although a discrepancy exists, experimental results can be reasonably well fitted by the hydrodynamic model [20] with a viscosity $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$, both showing a maximum friction when the spherical particle is symmetrically immersed, with its equator coinciding with the membrane plane. However, some data cannot be fitted with this value of $\eta_{m}$ and correspond to particles interacting with vesicles containing substructures, which may increase the viscosity of the membrane.
For fluid bilayers, the molecular rearrangement of lipids prevents the generation of shear stress across the membrane, thus the latter can be described as a thin region of liquid with viscosity $\eta$ and thickness $\delta$, without generation of velocity gradient, or equivalently, can be treated as an anisotropic material with a constant surface shear viscosity $\eta_{m}=\delta\eta$. Therefore, the increase in friction between particles and GUVs containing substructures in their vicinity can be attributed to the linear dependence of surface shear viscosity on membrane thickness, and hence on the number of lipid monolayers. Consequently, the two highest values of the friction (pink circles), deviating from the trend ($\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$), can be described by a (2.5 times) higher effective viscosity, possibly indicating the presence of a second bilayer.
Relating our findings to previous studies, Shigyou et al. [12] studied the diffusion of submicron polystyrene particles wrapped by DOPC GUVs (by centrifugation) and classified the wrapping states into two categories: ”full” or ”partial”. Due to the particle size (200 - 800 nm in diameter), intermediate wrapping states could not be optically resolved. They reported a higher drag for fully wrapped particles compared to partially wrapped ones, hence, by comparing with our data (although involving larger particles), this suggests that the partially wrapped particles in their study likely had very shallow penetration depths. In order to estimate membrane viscosity, their data were compared with the DADL model [18] [11], which assumes a flat membrane crossing the particle equator and does not account for local deformations around the particle or the presence of a water gap. Although this model is a simplified approach, they determined a membrane viscosity of $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$, the same value we found in our experiments. Recently, Van der Wel et al. [6] studied the adhesion between avidin-coated particles and vesicles containing biotinylated lipids, and observed two equilibrium states: fully wrapped or attached. Surprisingly, they reported higher diffusivity for fully wrapped particles, and a possible explanation for this result was qualitatively attributed to an increased membrane-particle distance compared to the attached state. However, these findings contrast with our results, where fully wrapped particles ($d/a=-2$) exhibit slower diffusion (higher friction) compared to attached particles ($d/a=0$) (see Fig. 4). In their study, the DADL model [18] was also applied to extract the membrane viscosity, yielding a value of $\eta_{m}=1.2\cdot 10^{-9}\ \mathrm{Pa.s.m}$. The impact of the particle-membrane distance, discussed in [6], which may be described as a water gap, is quantitatively analyzed in the following section.
@@ II.2.2 Simulations
Figure 6: Translational friction of a spherical particle (radius $a=0.6\,\mathrm{\mu m}$) as a function of penetration depth through a membrane, modeled as a layer of incompressible fluid (width $\delta=5\,\mathrm{nm}$) and constant surface shear viscosity $\eta_{m}$, separated by two identical regions of incompressible liquid (water). A water gap (thickness $l$) between the particle and the membrane is included. Our numerical results are compared to Fischer et al. calculations for surface shear viscosites $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$ and $\eta_{m}=10^{-8}\ \mathrm{Pa.s.m}$ (blue dotted and dashed curves). The points from the simulation, represented as circles and diamonds, are connected by a solid curve serving as a visual guide and does not represent an interpolation.
@@
Several studies have demonstrated the existence of a variable water gap thickness in solid surface-membrane systems. For instance, Rädler et al. [29] investigated the weak interaction of giant vesicles with flat rigid substrates and found a mean separation distance of about 40 nm. Other studies have reported the existence of a 1-3 nm water gap in supported bilayer systems [30] [31] (given the strong adhesion in our system, the particle-membrane equilibrium distance is likely near the lower bound of this range). Concerning viscosity, several authors have measured values for DOPC bilayers ranging from 1.2 to 15 $\cdot 10^{-9}\ \mathrm{Pa.s.m}$ [6] [12] [32] [33]. Fig. 6 shows the calculated values of the translational friction coefficient as a function of penetration depth for different water gap thicknesses and membrane viscosities, reflecting the range of data obtained in these different studies, with a water gap of 5 – 10 nm and viscosity ranging from 2 to 10 $\cdot 10^{-9}\ \mathrm{Pa.s.m}$.
Our simulations exhibit a trend similar to Fischer et al. calculations, with maximum friction at half-engulfment. However here, the deformed shape of the membrane induced by particle wrapping is taken into account and the evolution of friction is no longer symmetrical with respect to the position $d/a=-1$. These simulations provide a correction to the hydrodynamic model [20], allowing a more accurate description of our experimental results, also displaying asymmetry. In our simulations, the presence of a water gap separating the particle and the membrane was also taken into account. A reduction in the thickness of this water layer results in increased friction, more pronounced when the membrane is highly viscous. This variation is minor for a viscosity $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$ but becomes significant for a viscosity $\eta_{m}=10^{-8}\ \mathrm{Pa.s.m}$. In our experiments, with a lipid bilayer viscosity estimated at $\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$, the thickness of the water gap has thus a small effect on friction. Hence, dissipation primarily occurs in the regions close to the junction between the unbound membrane plane and the cap (formed by the particle portion engulfed and covered by the membrane, see Fig. 5A).
@@
Figure 7: (A) Schematic representations and bright-field images of a particle wrapped by a vesicle bilayer and a particle at the interface between the internal aqueous phase and the external oil phase of a droplet, stabilized by a lipid monolayer. The lipid composition used for the fabrication of the vesicles and droplets is identical. (B) Typical two-dimensional projected (in the (x,y) plane) particles trajectories in the two configurations, for an acquisition time of 10 s and vesicle/droplet radii ranging from 3.5 to 7 µm. (C) Mean squared displacement (calculated from the 3D reconstruction of the 2D projected trajectory of the particle) as a function of lag time for particles wrapped by a lipid bilayer, with different penetration depths and (D) particles wetting a lipid monolayer. In the case of particles encapsulated in droplets, the MSD scales linearly with time, even for longer timescales ($\Delta t\gg 0.03$ s), characteristic of purely Brownian motion, but exhibits a lower diffusion coefficient compared to particles wrapped by a bilayer (see insets). In this latter case, the transport is hindered and manifests as a sublinear increase of the MSD at longer timescales ($\mathrm{MSD}\propto\Delta t^{\alpha}$, with $\alpha=0.78\pm 0.14$ based on all the data analyzed). The diffusion is systematically lower in both cases compared to the Stokes-Einstein predictions for a sphere in the bulk (red curve).
@@ III Particle diffusion at the water/lipid monolayer/oil interface
The analysis of the diffusion of particles wrapped by GUVs, potentially containing substructures near their contact zone, indicates an influence of the effective membrane thickness on translational friction. To investigate further the effect of the layer thickness, we performed experiments using lipid monolayers showing a thickness that is half that of lipid bilayers. We encapsulated particles in water droplets stabilized by a lipid monolayer (same composition as the vesicles) and dispersed in oil (dodecane, $\eta_{o}=1.324\cdot 10^{-3}\ \mathrm{Pa.s}$ [34]).
Unlike GUVs, where particles spontaneously adhere to the outer leaflet of the lipid bilayer during the observation phase, here, the particles are directly encapsulated during the fabrication process. Fig. 7A illustrates the two systems: at the bottom, a particle partially engulfed by a vesicle bilayer and at the top, a particle encapsulated in a droplet, in close proximity with the monolayer.
In pure water, the particle does not adhere to the interface and diffuses freely in three dimensions within the aqueous droplet. The addition of NaCl in the aqueous phase (10 mM) results in a reduction of the Debye length and causes the particle to approach the interface. In this configuration, the particle does not deform the lipid monolayer but is in a state of partial wetting, moving along the curvature of the droplet and exploring a larger region than particles wrapped by a GUV bilayer. Fig. 7B depicts the projected 2D trajectories of particles encapsulated in droplets and wetting the interface (green) and of particles partially engulfed by GUVs (purple), over the same acquisition time (10 s), for vesicle/droplet radii ranging from 3.5 to 7 µm, clearly showing this phenomenon.
@@
Although the trajectories of particles encapsulated in droplets are more extensive, the mean squared displacement evaluated at short times (MSD linear with the lag time for $\Delta t<0.03$ s) reveals slower diffusion compared to particles wrapped by lipid bilayers. MSD analysis for particle-bilayer systems at longer times reveals a sublinear behaviour characteristic of confined diffusion dynamics (see Fig. 7C and D). The crossing from the diffusive and subdiffusive regimes occurs around $\Delta t\simeq 0.05$ s, and the subdiffusion is characterized with an exponent $\alpha=0.78\pm 0.14$ (no correlation was observed between $d/a$ and $\alpha$). This phenomenon is ubiquitous in cellular biology and has also been studied using model systems. Several works have shown that the presence of substructures (buds) [12] and membrane crowding induced by proteins [35] or lipid domains [36] lead to anomalous transport. In our system, the vesicles are made of only two types of lipids in the liquid phase (and 1% of a fluorescently labeled lipid), and thus diffusion cannot be perturbed by complex systems as proteins. The presence of substructures may explain this type of long time diffusion, but in our system, it appears even in GUVs without optically visible substructures. The finite-size effects induced by the spherical shape of the vesicle are not the cause of this confinement either, as it does not occur for a particle encapsulated in a droplet of the same size. Another hypothesis explaining particle subdiffusion could be the formation of an electrostatically induced DOTAP-rich zone in its vincinity. Among the various factors possibly contributing to particle subdiffusion, the role of membrane elasticity has also been investigated. Daddi-Moussa-Ider et al. analytically demonstrated that the motion of a particle near a membrane causes its deformation, followed by relaxation, generating a flow acting back on the particle with a certain delay. This memory effect, induced by the elastic nature of the membranes, is responsible for a transient subdiffusive phenomenon [37]. Although their work describes the diffusion of a particle near a membrane rather than being wrapped, this interpretation remains compelling in explaining the experimentally observed anomalous diffusion.
Going back to the short lag time diffusive regime, we measured a particle diffusion at the lipid monolayer interface significantly lower than for lipid bilayers even if the thickness of the monolayer is half that of the bilayer. Using the hydrodynamic model [20] to fit our experimental results for monolayers yields a surface shear viscosity of approximately $1.5\cdot 10^{-8}\ \mathrm{Pa.s.m}$, which is substantially higher than the estimated viscosity of lipid bilayers, even those containing substructures (in the range $\eta_{m}=2\cdot 10^{-9}-5\cdot 10^{-9}\ \mathrm{Pa.s.m}$, see Fig. 4).
@@
In this model and our simulations, the interface is described as flat (no roughness), since interfacial tension stresses dominate over viscous stresses, with a capillary number $Ca=\eta U/\sigma_{wo}\ll 1$ where $U$ is the particle velocity and $\sigma_{wo}$ is the water-oil interfacial tension. The triple-phase contact line between the particle and the two immiscible liquids appears stationary at the microscopic scale, but at the molecular level, it undergoes significant thermal motion, leading to additional energy dissipation. To more precisely capture this behavior, we used the approach of Boniello et al. [38] where a spherical microparticle diffuses at the water/air interface with a contact angle $\theta$ (related to the penetration depth by $1+\cos\theta=-d/a$), satisfying the local equilibrium of the triple contact line. Thermally activated displacements of the interface at the contact line induce additional random forces acting on the particle, which are associated with an additional viscous friction through the fluctuation-dissipation theorem. The triple-phase contact line fluctuates around its average position with a characteristic deformation length scale denoted as $\lambda$, inducing a line friction acting on the particle surface $\gamma_{tr,L}$ defined by:
@@
with $\chi=\arctan({4\sin\theta/\pi})$ being the angle between the horizontal plane and the tangents of the interface on the particle surface and $\tau=\kappa_{0}^{-1}$ corresponding to the correlation time of the interfacial fluctuations. This correlation time differs from that applied in the article by Boniello et al. describing a water/air interface. Here, both media are liquids with viscosities $\eta_{w}$ (water) and $\eta_{o}$ (oil) and the equilibrium frequency of random molecular displacements, $\kappa^{0}$, is defined by [39]:
@@
where $v_{w}=0.03\ \mathrm{nm}^{3}$ and $v_{o}=0.38\ \mathrm{nm}^{3}$ denote the molecular volume of water and oil, $h$ the Planck constant and $\sigma_{wo}(1+\cos\theta)$ corresponds to the reversible work of adhesion between the solid and the liquid, while considering the presence of the other liquid phase. Fig. 8 schematically represents the particle diffusing at the water/lipid monolayer/oil interface and the various parameters involved.
The total friction experienced by the particle diffusing at the lipid monolayer interface is equal to the sum of two uncorrelated terms $\gamma_{tr}=\gamma_{tr,H}+\gamma_{tr,L}$. For the calculation of the hydrodynamic friction $\gamma_{tr,H}$, we approximate the system as a particle dragged along a liquid-liquid interface, with both liquids having identical viscosities, using the result of the hydrodynamic model in [20], the ratio $\eta_{o}/\eta_{w}=1.324$ being close to 1. We set the surface viscosity of the monolayer to $\eta_{m}/2$, where $\eta_{m}$ is the viscosity of a membrane bilayer with the same lipid composition ($\eta_{m}=2\cdot 10^{-9}\ \mathrm{Pa.s.m}$, see Fig. 4). The interfacial tension for similar systems was measured using the pendant drop technique and found to vary between $2\text{-}5\,\mathrm{mN.m^{-1}}$ [40] [41]. Considering a small contact angle $5^{\circ}<\theta<10^{\circ}$, the characteristic deformation length $\lambda$ ranges between 2.7 and 4.8 nm which can be related to thermally excited capillary waves for low tension interfaces [42] and is comparable to the length of a lipid molecule.
This analysis reveals a fundamental difference between the diffusion mechanisms of a particle wrapped by a lipid bilayer, well described by a hydrodynamic model, and a particle diffusing near the water/lipid monolayer/oil interface of a droplet. For the latter system, fluctuations of the three-phase contact line must be considered, as they introduce an additional contribution to the translational drag experienced by the particle diffusing at the interface. The result of our analysis can be used to explain experiments reported in the literature on the slow lateral diffusion of nano- and micro-particles at the oil-water interface, which could not be rationalised so far [43].